OpenStudy (anonymous):

(5+a)squaredplusb a=2 b=5

OpenStudy (anonymous):

5+2=7^2 =49 49+5=54

OpenStudy (anonymous):

54 is your answer

OpenStudy (anonymous):

or is it (5+a)^(2+b) ??